OpenStudy (anonymous):

How many radians on the unit circle would the minute hand travel from 0° if it were to move 3π inches? What is the coordinate point associated with this radian measure?

OpenStudy (anonymous):

@satellite73 @Loser66

OpenStudy (anonymous):

I'm not sure if this is needed but here it is, Freddie is at chess practice waiting on his opponent's next move. He notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle.

OpenStudy (anonymous):

@rational

OpenStudy (rational):

Is the length of minute hand 4 in ?

OpenStudy (anonymous):

Yes

OpenStudy (rational):

The length travelled on "unit circle" equals the angle in radian measure. So simply find the angle intercepted in travelling 3pi inches on the clock

OpenStudy (rational):

\[l=r\theta\] plugin \(l=3\pi\), \(r=4\) and solve \(\theta\)

OpenStudy (anonymous):

is it just \[\frac{ 3\pi }{ 4 }\]

OpenStudy (rational):

Yep thats the angle in radians also, thats the distance travelled on unit circle

OpenStudy (rational):

that doesn't make sense could you show it in more steps

OpenStudy (rational):

do you mean l = r(theta) 3pi = 4(theta) 3pi / 4 = theta ?

OpenStudy (rational):

if so, looks good.

OpenStudy (anonymous):

yes haha

OpenStudy (rational):

just so you know, it is important to show the steps correctly. there is no way for others to predict what you think/assume

OpenStudy (anonymous):

ok, what about the second part of the question for coordinate plane?

OpenStudy (anonymous):

coordinate point

OpenStudy (rational):

A point has two coordinates : x, y

OpenStudy (rational):

you can find them using : \(x = r\cos(\theta)\) \(y=r\sin(\theta)\)

OpenStudy (rational):

\(r\) = radius of clock = \(4\) \(y\) = angle from 0 degrees = \(\frac{3\pi}{4}\) plug them

OpenStudy (anonymous):

just \[4\cos (\theta)\] ?

OpenStudy (rational):

\(x = r\cos(\theta) = 4*\cos(\frac{3\pi}{4}) = ?\)

OpenStudy (anonymous):

i get \[\sqrt[-2]{2}\] what does that simplify into?

OpenStudy (rational):

you should be getting \(\large x = -2*\sqrt{2}\)

OpenStudy (rational):

and \(\large y = 2*\sqrt{2}\) so the required coordinate point would be \[\large (-2\sqrt{2},~2\sqrt{2})\]

OpenStudy (anonymous):

okay thank you, that's a weird coordinate,

OpenStudy (rational):

np:)

OpenStudy (rational):

btw, \(\sqrt{2}\) is just a number